# Dust evolution during the protostellar collapse: influence on the coupling between the neutral gas and the magnetic field

### v1

@@ 6.3 Uncertainties in the turbulence model
In addition, turbulence can affect the local dust-to-gas ratio at small scales which is an essential ingredient in the context of dust growth. Indeed the turbulent coagulation timescale scales as the inverse of the dust-to-gas ratio squared. Turbulent motions are known to be capable of triggering hydrodynamics instabilities, allowing the dust-to-gas ratio to increase in certain regions (Kelvin-Helmholtz instability for instance, see Hendrix & Keppens, 2014)). Dust clumping by supersonic turbulence has been observed numerically in an environment reproducing molecular cloud physical conditions in Commerçon et al. (2023), and in Hopkins & Lee (2016) for supersonic MHD turbulence. With both an analytical and numerical approach, Hopkins & Squire (2018b) expanded their own work on the Resonant Drag Instability (RDI) by including the role of magnetic fields and charged dust. As the dust and the gas drift relatively to one another (via gravity for example), instabilities develop for all modes wavelength, leading to strong fluctuations in the dynamics of both fluids. In addition to the purely hydrodynamical instabilities observed in their previous work (Hopkins & Squire, 2018a; Squire & Hopkins, 2018), they identified a whole set of new instabilities. More generally, hydrodynamical instabilities represent promising mechanisms to allow the formation of planetesimals.
@@ 7 Conclusion
In this article, we studied how the evolution of the dust distribution could affect the degree of coupling between the gas and the magnetic field of a collapsing core, including both coagulation and fragmentation of dust grains. We also investigated the effect of grain-grain erosion but neglected other collision outcomes such as compaction, bouncing and rotational disruption. We used the shark code to follow self-consistently the evolution of the dust distribution and the subsequent impact on the magnetic resistivities. The code includes brownian motion, turbulence and ambipolar drift as grain-grain collision sources. We performed single zone simulations to save computational time and focus on the influence of fragmentation and various parameters.
To summarize, we found that :
@@
    Likewise previous studies (Guillet et al., 2020; Kawasaki et al., 2022; Lebreuilly et al., 2023; Marchand et al., 2023), dust growth is found to be significant and the subsequent impact on the magnetic resistivities very important. Therefore, dust evolution cannot be disregarded in numerical simulations.
@@
    The degree of influence of the parameters depends on the fragmentation efficiency, which is primarily controlled by the grain surface energy. When considering a high value for this parameter ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$, i.e., resistant grains), fragmentation is substantially suppressed and the magnetic resistivities are only sensitive to the parameters which do not affect only (or not at all) the fragmentation processes, namely the ambipolar drift intensity, the grain-grain erosion efficiency, electrostatic repulsion and the cosmic ray ionization rate. However when considering a low value for the grain surface energy ($\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, fragile grains), fragmentation is severe and fragmentation-related parameters such as the monomer size, the grain intrinsic density and the power-law index of the fragment redistribution law become influential.
@@
    Grain fragmentation offers a way to replenish the population of small grains, which in turn significantly affects the resistivities. However, except for a very severe fragmentation, we find that the small grains are very rapidly removed due to ambipolar drift. As a consequence, the ohmic resistivity $\eta_{\mathrm{O}}$ remains negligible with respect to the ambipolar resistivity $\eta_{\mathrm{AD}}$ for the entirety of the collapse. In addition, the ambipolar resistivity is much higher at low densities and much lower at high densities than the value one gets in considering a fixed MRN dust distribution. Increasing the initial magnetic field strength leads to a more efficient depletion of small grains by ambipolar drift and a sharper rise in the ambipolar resistivity. The magnetic field then diffuses more rapidly, thus hinting at a regulation of the magnetic flux by ambipolar diffusion.
@@
    To alleviate the tension between the remarkably high ambipolar resistivities obtained in our simulations and the statistically low protoplanetary disk sizes observed, we identified a few mechanisms that allow to reduce the ambipolar resistivity $\eta_{\mathrm{AD}}$ and thus supposedly enhance magnetic braking : low grain surface energy (i.e., fragile grains), large ionization rate by cosmic-rays and low grain-grain sticking efficiency. In addition, we showed that for a sufficiently low ambipolar drift on one hand, grain-grain electrostatic repulsion efficiently inhibits small grain depletion, leading to a much lower ambipolar resistivity prior to $n_{\mathrm{H}}=10^{11}\ \mathrm{cm^{-3}}$ (whereas grain-grain erosion is uninfluential). On the other hand in the case of a strong ambipolar drift, while the Coulomb barrier for electrostatic repulsion is easily overcome, grain-grain erosion causes a significant replenishment of the small grains and consequently a sharp reduction of the ambipolar resistivity. Both effects are expected to maintain a reasonable ambipolar resistivity $\eta_{\mathrm{AD}}$ during the protostellar collapse.
@@
    We pointed out the uncertainties regarding the elastic properties of the grains and the fragmentation model. Given the dramatic influence fragmentation can have on the magnetic resistivities, it is fundamental to better constrain the grain surface energy and the various collision outcomes relevant to interstellar dust in the adequate range of grain sizes. Besides, the grain surface energy also controls the maximum size dust grains can reach by the end of the collapse. For $\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, the grains grow up to $0.1\ \mathrm{mm}$ while for $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$, we manage to form grains of a few $\mathrm{mm}$.
As a last note, we recall that we used the square root approximation of Eq. (21) to compute the evolution of the magnetic field during the gravitational collapse as a function of the gas density only. However, the magnetic field affects the resistivities, which should in turn control the evolution of the latter through the induction equation, so on so forth. To rigorously describe the evolution of the magnetic field, multidimensional simulations including dust growth along with a self-consistent computation of the magnetic field are needed.
@@ Acknowledgements.
Appendix A What species controls the conductivities during the collapse?
@@
Figure 8 explicitly shows that the individual electronic conductivity largely prevails over the others by a few orders of magnitude. Consequently, the total parallel conductivity is controlled by the electrons. In reducing the surface energy of the grains $\gamma_{\mathrm{grain}}$, we allow the small grains to be repopulated and to efficiently collect the electrons, resulting in a weaker overall parallel conductivity. Since the collision rate between small grains and electrons increases with density (regardless of dust coagulation), the impact on the electron conductivity and thus on the parallel conductivity only appears tardily in the collapse. Indeed, the dashed line diverges from the solid line at high densities only. The same reasoning is valid for the ions whose conductivity decreases as they recombine with electrons at the grains surface.
Regarding the individual perpendicular conductivities, the electronic contribution is negligible. Indeed, Fig. 9 suggests that the total perpendicular conductivity is dominated by both the dust and the ions at low density. Eventually as the density goes up, the dust influence vanishes and the ions prevail. In enhancing fragmentation, the replenishment of small grains gives rise to a higher dust perpendicular conductivity (the large grains contribution is insignificant) which leads to a higher total perpendicular conductivity at low density ($n_{\mathrm{H}}\leq 10^{9}\ \mathrm{cm^{-3}}$). This effect is instantaneous since it does not involve any capture or interaction whatsoever. On the third panel, we can see that the dashed and solid lines overlap as soon as the dust ceases to be the dominant species, and the detachment between both curves at high densities is due to the ion density diminution (caused by electron/ion recombination) being significant at high densities only. We can derive an analytical formulae of a threshold value above which the ion density allows the ion perpendicular conductivity to be larger than the dust perpendicular conductivity, i.e., when $\sigma_{\mathrm{perp,ion}}>\sigma_{\mathrm{perp,dust}}$. When this condition is satisfied, the ambipolar resistivity $\eta_{\mathrm{AD}}$ becomes independent on the number of small grains, which explains why strong variations in the population of small grains sometimes do not reflect on the ambipolar resistivity (see for example Fig. 3 at $n_{\mathrm{H}}=10^{8}\ \mathrm{cm^{-3}}$). Using the expressions of Eq. (3.1.4), we get
@@
where $m_{\mathrm{d}}$, $\tau_{\mathrm{d}}$, $n_{\mathrm{ion}}$, $\mu_{\mathrm{ion}}$ and $\tau_{\mathrm{ion}}$ are respectively the dust mass and reduced temperature, the ion density, mean ion molecular weight and reduced temperature. $\frac{dN_{\mathrm{d}}}{ds}=As^{-q}$ is the dust number distribution taken as a power law of index $q$ and $A=\frac{(4-q)\rho_{\mathrm{dust}}}{\frac{4}{3\pi}\rho_{\mathrm{grain}}\mu_{\mathrm{gas}}|s_{\mathrm{max}}^{4-q}-s_{\mathrm{min}}^{4-q}|}$ is a normalisation factor with $s_{\mathrm{max}}$ and $s_{\mathrm{min}}$ being the uppermost and lowermost sizes of the dust distribution considered. Integrating for any $q<4$, we get
@@
$K\left(n_{\mathrm{H}},T\right)=\frac{\epsilon m_{\mathrm{H}}\pi^{2}\mu_{\mathrm{gas}}n_{\mathrm{H}}c_{\mathrm{s}}\tau_{\mathrm{ion}}}{\rho_{\mathrm{grain}}\mu_{\mathrm{ion}}\sqrt{\frac{\pi\gamma}{8}}}$ is a factor which depends on the gas density and the temperature. Considering that $\frac{s_{\mathrm{min}}}{s_{\mathrm{max}}}\ll 1$, we can simplify the expressions
@@
We can see that in reducing $s_{\mathrm{min}}$, the dust contribution would be more important and thus the ion density would have to be higher in order to maintain the dominance of the ion contribution. In increasing $s_{\mathrm{max}}$, for a given $q$ and $s_{\mathrm{min}}$, we reduce the number of small grains and consequently relax the constrain on $n_{\mathrm{ion}}$ which can afford to be lower. We stress that to apply this equation at some density $n_{\mathrm{H}}$, one has to describe precisely the dust distribution, that is to know $s_{\mathrm{max}}$, $s_{\mathrm{min}}$ and $q$. To gain insight, we apply Eq. (31) at the initial conditions of the collapse for which the shape of the distribution is known, since it is set as an MRN-like power law. With $s_{\mathrm{min}}=5\ \mathrm{nm}$, $s_{\mathrm{max}}=250\ \mathrm{nm}$, $q=3.5$, $n_{\mathrm{H}}=2.6\times 10^{6}\ \mathrm{cm^{-3}}$ and $T=10\ \mathrm{K}$, we get $n_{\mathrm{ion}}>1.67\ \mathrm{cm^{-3}}$, and we can see on the various plots that initially $n_{\mathrm{ion}}\thicksim 10^{-2}\ \mathrm{cm^{-3}}$, which means it is indeed the dust contribution which prevails in regard to the perpendicular conductivity at that particular moment.
@@ Appendix B Comparison between fixed MRN and coagulating distribution with and without fragmentation
Appendix C What is is the most efficient source of relative velocity for the removal of small grains ?
@@
Figure 13 provides a comparison of the effect of the different sources of grain collisions on the removal of small grains. $\mathrm{TURB-COAG}$ model includes only turbulence as a source of grain collisions. $\mathrm{TURB-BR-COAG}$ includes additionally Brownian motion while $\mathrm{TURB-AD-COAG}$ contains ambipolar diffusion. The models displaying $\mathrm{FRAG}$ account for fragmentation in addition to coagulation. We see that in the case of pure coagulation, the number of small grains is much lower when ambipolar diffusion is included, while it remains higher when only brownian motion is accounted for, at all densities. Therefore in this context, ambipolar diffusion is the main process in regard to the depletion of small grains. However when including fragmentation, it seems that the gap in small grain removal efficiency between ambipolar diffusion and brownian motion is clearly reduced as fragmentation allows for a systematic replenishment of the small grains. At intermediate densities, ambipolar diffusion still appears to be slightly more efficient at removing the small grains, whereas at high density (last column) both velocity sources are contributing to the depletion of small grains with Brownian motion being slightly more efficient. Note that since $\mathrm{TURB-AD-FRAG}$, $\mathrm{TURB-BR-FRAG}$ and $\mathrm{TURB-FRAG}$ curves are not completely overlapping, small grain depletion is not completely suppressed by the repopulation induced by fragmentation. In addition, although there exists a striking difference in small grain population between models that include ambipolar diffusion and those which do not, we still observe a reduction of a few orders of magnitude in the overall number of small grains during the collapse (even more striking when fragmentation is off) regardless of the velocity source at play. Consequently, we can see on Fig. 14 that even without ambipolar drift, the ambipolar resistivity $\eta_{\mathrm{AD}}$ reaches values larger than the MRN case.

### v2

@@ 6.3 Uncertainties in the turbulence model
In addition, turbulence can affect the local dust-to-gas ratio at small scales which is an essential ingredient in the context of dust growth. Indeed the turbulent coagulation timescale scales as the inverse of the dust-to-gas ratio squared. Turbulent motions are known to be capable of triggering hydrodynamics instabilities, allowing the dust-to-gas ratio to increase in certain regions (Kelvin-Helmholtz instability for instance, see Hendrix & Keppens, 2014)). Dust clumping by supersonic turbulence has been observed numerically in an environment reproducing molecular cloud physical conditions in Commerçon et al. (2023), and in Hopkins & Lee (2016) for supersonic MHD turbulence. With both an analytical and numerical approach, Hopkins & Squire (2018b) expanded their own work on the Resonant Drag Instability (RDI) by including the role of magnetic fields and charged dust. As the dust and the gas drift relative to each other (e.g. via gravity), instabilities develop for all modes wavelength, leading to strong fluctuations in the dynamics of both fluids. In addition to the purely hydrodynamical instabilities observed in their previous work (Hopkins & Squire, 2018a; Squire & Hopkins, 2018), they identified a whole set of new instabilities. More generally, hydrodynamical instabilities represent promising mechanisms to enable the formation of planetesimals.
@@ 7 Conclusion
In this work, we study how the evolution of the dust distribution could affect the degree of coupling between the gas and the magnetic field of a collapsing core, including both the coagulation and fragmentation of dust grains. We also investigated the effect of grain-grain erosion, but we neglected other collision outcomes such as compaction, bouncing, and rotational disruption. We used the shark code to self-consistently follow the evolution of the dust distribution and the subsequent impact on the magnetic resistivities. The code includes brownian motion, turbulence and ambipolar drift as grain-grain collision sources. We performed single zone simulations to save computational time and focus on the influence of fragmentation and various parameters. In summary, we found:
@@
    As in previous studies (Guillet et al., 2020; Kawasaki et al., 2022; Lebreuilly et al., 2023; Marchand et al., 2023), dust growth is found to be significant and the subsequent impact on the magnetic resistivities very important. Therefore, dust evolution cannot be disregarded in numerical simulations.
@@
    The degree of influence of the parameters depends on the fragmentation efficiency, which is primarily controlled by the grain surface energy. When considering a high value for this parameter ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$, i.e. resistant grains), fragmentation is substantially suppressed and the magnetic resistivities are only sensitive to the parameters which do not affect only (or not at all) the fragmentation processes; namely, the ambipolar drift intensity, the grain-grain erosion efficiency, electrostatic repulsion, and the cosmic-ray ionisation rate. However, when considering a low value for the grain surface energy ($\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, fragile grains), fragmentation is severe and fragmentation-related parameters, such as the monomer size, the grain intrinsic density, and the power-law index of the fragment redistribution law become influential.
@@
    Grain fragmentation offers a way to replenish the population of small grains, which, in turn, significantly affects the resistivities. However, except for cases of very severe fragmentation, we find that the small grains are very rapidly removed due to ambipolar drift. As a consequence, the ohmic resistivity, $\eta_{\mathrm{O}}$, remains negligible with respect to the ambipolar resistivity, $\eta_{\mathrm{AD}}$, for the entirety of the collapse. In addition, the ambipolar resistivity is much higher at low densities and much lower at high densities than the value one gets in considering a fixed MRN dust distribution. Increasing the initial magnetic field strength leads to a more efficient depletion of small grains by ambipolar drift and a sharper rise in the ambipolar resistivity. The magnetic field then diffuses more rapidly, thus hinting at the regulation of the magnetic flux by ambipolar diffusion.
@@
    To alleviate the tension between the remarkably high ambipolar resistivities obtained in our simulations and the statistically low protoplanetary disc sizes observed, we identified a few mechanisms that allow for the ambipolar resistivity, $\eta_{\mathrm{AD}}$, to be reduced and, thus, to supposedly enhance magnetic braking: low grain surface energy (i.e. fragile grains), high ionisation rate via cosmic rays and low grain-grain sticking efficiency. In addition, we show that for a sufficiently low ambipolar drift on one hand, grain-grain electrostatic repulsion efficiently inhibits small grain depletion, leading to a much lower ambipolar resistivity prior of $n_{\mathrm{H}}=10^{11}\ \mathrm{cm^{-3}}$ (whereas the grain-grain erosion is negligible). On the other hand, in the case of a strong ambipolar drift, the Coulomb barrier for electrostatic repulsion is easily overcome, however, grain-grain erosion causes a significant replenishment of the small grains; consequently, there is a sharp reduction of the ambipolar resistivity as a result. Both effects are expected to maintain a reasonable ambipolar resistivity, $\eta_{\mathrm{AD}}$, during the protostellar collapse.
@@
    We pointed out the uncertainties regarding the elastic properties of the grains and the fragmentation model. Given the dramatic influence fragmentation can have on the magnetic resistivities, it is fundamental to better constrain the grain surface energy and the various collision outcomes relevant to interstellar dust in the adequate range of grain sizes. In addition, the grain surface energy also controls the maximum size dust grains can reach by the end of the collapse. For $\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, the grains grow up to $0.1\ \mathrm{mm}$ while for $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$, we manage to form grains of a few $\mathrm{mm}$.
As a last note, we recall that we used the square root approximation of Eq. (21) to compute the evolution of the magnetic field during the gravitational collapse as a function of the gas density only. However, the magnetic field affects the resistivities, which should (in turn) control the evolution of the latter through the induction equation and so forth. To rigorously describe the evolution of the magnetic field, multidimensional simulations, including dust growth along with a self-consistent computation of the magnetic field are needed.
@@ Acknowledgements.
Appendix A Which species controls the conductivities during collapse
@@
Figure 8 explicitly shows that the individual electronic conductivity largely prevails over the others by a few orders of magnitude. Consequently, the total parallel conductivity is controlled by the electrons. In reducing the surface energy of the grains, $\gamma_{\mathrm{grain}}$, we allow for the small grains to be repopulated and to efficiently collect the electrons, resulting in a weaker overall parallel conductivity. Since the collision rate between small grains and electrons increases with density (regardless of dust coagulation), the impact on the electron conductivity and thus on the parallel conductivity only appears tardily in the collapse. Indeed, the dashed line diverges from the solid line at high densities only. The same reasoning is valid for the ions whose conductivity decreases as they recombine with electrons at the grains surface.
Regarding the individual perpendicular conductivities, the electronic contribution is negligible. Indeed, Fig. 9 suggests that the total perpendicular conductivity is dominated by both the dust and the ions at low density. Eventually as the density goes up, the dust influence vanishes and the ions prevail. In enhancing fragmentation, the replenishment of small grains gives rise to a higher dust perpendicular conductivity (the large grains contribution is insignificant) which leads to a higher total perpendicular conductivity at low density ($n_{\mathrm{H}}\leq 10^{9}\ \mathrm{cm^{-3}}$). This effect is instantaneous since it does not involve any capture or interaction whatsoever. On the third panel, we can see that the dashed and solid lines overlap as soon as the dust ceases to be the dominant species, and the detachment between both curves at high densities is due to the ion density diminution (caused by electron/ion recombination) being significant at high densities only. We can derive an analytical formulae of a threshold value above which the ion density allows for the ion perpendicular conductivity to be larger than the dust perpendicular conductivity, namely, when $\sigma_{\mathrm{perp,ion}}>\sigma_{\mathrm{perp,dust}}$. When this condition is satisfied, the ambipolar resistivity $\eta_{\mathrm{AD}}$ becomes independent on the number of small grains, which explains why strong variations in the population of small grains sometimes do not reflect on the ambipolar resistivity (see for example Fig. 3 at $n_{\mathrm{H}}=10^{8}\ \mathrm{cm^{-3}}$). Using the expressions of Eq. (3.1.4), we get:
@@
where $m_{\mathrm{d}}$, $\tau_{\mathrm{d}}$, $n_{\mathrm{ion}}$, $\mu_{\mathrm{ion}}$ and $\tau_{\mathrm{ion}}$ are respectively the dust mass and reduced temperature, the ion density, mean ion molecular weight and reduced temperature. $\frac{dN_{\mathrm{d}}}{ds}=As^{-q}$ is the dust number distribution taken as a power law of index $q$ and $A=\frac{(4-q)\rho_{\mathrm{dust}}}{\frac{4}{3\pi}\rho_{\mathrm{grain}}\mu_{\mathrm{gas}}|s_{\mathrm{max}}^{4-q}-s_{\mathrm{min}}^{4-q}|}$ is a normalisation factor with $s_{\mathrm{max}}$ and $s_{\mathrm{min}}$ being the uppermost and lowermost sizes of the dust distribution considered. Integrating for any $q<4$, we get:
@@
Here, $K\left(n_{\mathrm{H}},T\right)=\frac{\epsilon m_{\mathrm{H}}\pi^{2}\mu_{\mathrm{gas}}n_{\mathrm{H}}c_{\mathrm{s}}\tau_{\mathrm{ion}}}{\rho_{\mathrm{grain}}\mu_{\mathrm{ion}}\sqrt{\frac{\pi\gamma}{8}}}$ is a factor which depends on the gas density and the temperature. Considering $\frac{s_{\mathrm{min}}}{s_{\mathrm{max}}}\ll 1$, we can simplify the expressions to:
@@
We can see that in reducing $s_{\mathrm{min}}$, the dust contribution would be more important and, thus, the ion density would have to be higher in order to maintain the dominance of the ion contribution. In increasing $s_{\mathrm{max}}$, for a given $q$ and $s_{\mathrm{min}}$, we reduce the number of small grains and consequently relax the constrain on $n_{\mathrm{ion}}$ which can afford to be lower. We stress that to apply this equation at some density $n_{\mathrm{H}}$, one has to describe precisely the dust distribution, that is to know $s_{\mathrm{max}}$, $s_{\mathrm{min}}$ and $q$. To gain insight, we apply Eq. (31) at the initial conditions of the collapse for which the shape of the distribution is known, since it is set as an MRN-like power law. With $s_{\mathrm{min}}=5\ \mathrm{nm}$, $s_{\mathrm{max}}=250\ \mathrm{nm}$, $q=3.5$, $n_{\mathrm{H}}=2.6\times 10^{6}\ \mathrm{cm^{-3}}$ and $T=10\ \mathrm{K}$, we get $n_{\mathrm{ion}}>1.67\ \mathrm{cm^{-3}}$, and we can see on the various plots that initially $n_{\mathrm{ion}}\thicksim 10^{-2}\ \mathrm{cm^{-3}}$, which means it is indeed the dust contribution which prevails in regard to the perpendicular conductivity at that particular moment.
@@ Appendix B Comparison between fixed MRN and coagulating distribution with and without fragmentation
Appendix C Considering the most efficient source of relative velocity for the removal of small grains
@@
Figure 13 provides a comparison of the effect of the different sources of grain collisions on the removal of small grains. $\mathrm{TURB-COAG}$ model includes only turbulence as a source of grain collisions. $\mathrm{TURB-BR-COAG}$ includes additionally Brownian motion while $\mathrm{TURB-AD-COAG}$ contains ambipolar diffusion. The models displaying $\mathrm{FRAG}$ account for fragmentation in addition to coagulation. We see that in the case of pure coagulation, the number of small grains is much lower when ambipolar diffusion is included, while it remains higher when only brownian motion is accounted for, at all densities. Therefore, in this context, ambipolar diffusion is the main process in regard to the depletion of small grains. However when including fragmentation, it seems that the gap in small grain removal efficiency between ambipolar diffusion and brownian motion is clearly reduced as fragmentation allows for a systematic replenishment of the small grains. At intermediate densities, ambipolar diffusion still appears to be slightly more efficient at removing the small grains, whereas at high density (last column), both velocity sources are contributing to the depletion of small grains with Brownian motion being slightly more efficient. Note that since $\mathrm{TURB-AD-FRAG}$, $\mathrm{TURB-BR-FRAG,}$ and $\mathrm{TURB-FRAG}$ curves are not completely overlapping, small grain depletion is not completely suppressed by the repopulation induced by fragmentation. In addition, although there exists a striking difference in small grain population between models that include ambipolar diffusion and those which do not, we still observe a reduction of a few orders of magnitude in the overall number of small grains during the collapse (even more striking when fragmentation is off) regardless of the velocity source at play. Consequently, we can see on Fig. 14 that even without ambipolar drift, the ambipolar resistivity, $\eta_{\mathrm{AD}}$, reaches values larger than the MRN case.